Systems Staff Meeting Notes 4/17/00

Systems Staff Meeting Notes
April 17, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Hilton Soto, Ying Tang
Announcements
  1. We'll be getting a new Programmer position. It will be located in Suzy's old cubicle. It is the DLC position that will move to Systems under Winston. It is an A&P coordinator position. We will be supporting the DLC interface programming, although the position will not be exclusive to DLC. The position has not been hired yet; the job description is not yet written. Developments on this position should take place this week.
  2. Everett's soccer team lost.
  3. Everett has intramural consolation softball finals on Wednesday.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. The rest of the FCLA money will be moving out - new purchases of machines for AC, E-Reserves and some new barcode readers will be forthcoming.
Networking
  1. IPX problems continue - John Ashcraft called last Thursday reporting CD access & all public printing was down, then called saying it was back up. Some changes were made on Sunday so that the ACL will get the nearest server (done through NERDC core server). This should improve access during the morning rush hour. Some rogue printers & one jet direct were found screaming at the IPX core router. This situation was corrected; hopefully this will help to resolve the problem. This will be particularly important when public printing expands to Journalism.
Current Projects
  1. NT 10 is ready - Will would like to run the test to see if it works on Friday (switching back to the Novell), giving him the time to troubleshoot over the weekend. The CD ROMs will be down for about two hours. Rich suggests doing it around 2 PM - he will check with the librarians and let Will know.
  2. Moving hubs to switches - fix will be submitted to Bill today.
  3. Wireless - on Brad's list to install the conduit from the stacks to the 1st floor Com closet. Bill suggests checking the end points to be sure it meets our needs. Brad will be joining to the existing conduit.
  4. All of the Dells have been distributed.
  5. Norton vs. McAfee conflict on Dells - new Dells do not have any anti-virus software installed.
  6. West wiring - still on hold for the cable tray guys. Digital should be able to resume wiring by the end of this week.
  7. Journalism - do we have sufficient floor mold for the job? No, but Bobbie informed Gary that is between HS&S and Facilities.
Software Projects Status
  1. Hilton and Winston will present the film log demo this week. Once the modifications are made he can start on LASSIE.
  2. Ying is still working on the Systems Wiring database. She needs user input.
  3. Ying is working to set up the data structure for task tracking on the server, in case it can be used via web interface.
DP Coordinator
  1. Everett will be managing SYSHELP/Trouble Tickets while Debra is gone for the rest of this week. Access to the SYSHELP mailbox has been added to the Systems group - you just need to add it to your mailbox file list to access it.
Calendar
  1. Today is tax day.
  2. Winston is out April 17-19.
  3. Debra is out April 18-21.
  4. Bobbie is out April 20-21.
  5. Rich is out April 24-28.
  6. Winston out June 19-23.
  7. Bobbie out July 10-14.
  8. Will out July 24-28.


Back to the Systems Homepage

Last updated April 24, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu